# Regularized Black Hole Solution from a New String Cloud Source

### v1

@@ Abstract
We construct a new family of regular black hole solutions supported by the novel Letelier–Alencar string cloud and regularized through a rational Dagum-type distribution. The regulator smooths the matter profile and ensures finite curvature invariants, yielding a geometry that interpolates between a string-cloud exterior and an anti–de Sitter core. We analyze the energy conditions, identifying where the null, weak, dominant and strong conditions hold or fail across the core and exterior. The parameter space for horizon formation is mapped and the thermodynamic propertie – mass, Hawking temperature, entropy and heat capacity – are derived; notably, the entropy depends only on the regularization scale while the string parameter modifies temperature and heat capacity. Employing Rényi non-extensive entropy and the topological thermodynamics approach, we show the non-extensive deformation stabilizes the system and removes the standard phase transition. Finally, we compute the shadow radius and derive constraints compatible with current Event Horizon Telescope bounds for Sgr A* and M87*.
@@ I Introduction
The search for regular solutions for black holes has become one of the most important topics in recent literature. Since the formulation of the Schwarzschild and Kerr metrics, the presence of central singularities (regions where curvature invariants diverge and General Relativity ceases to be valid) has been interpreted as an indication of the classical incompleteness of the theory. Although the singularity theorems developed by Penrose and Hawking indicate that such divergences are inevitable under physically reasonable energy conditions[hawking1971gravitational, wald2001thermodynamics], it is widely believed that the fundamental theory of gravitation should remain regular in high curvature regimes, in which quantum effects or nonlinear corrections become relevant [rovelli2004quantum].
@@
In this context, the construction of regular black holes (RBNs) offers a particularly useful alternative to eliminate singularities, introducing smooth energy and momentum distributions that produce finite curvature invariants throughout spacetime. Early models, such as those proposed by [Bardeen1973, dymnikova1992vacuum, hayward2006formation], replaced the central singularity with a de Sitter nucleus (dS) characterized by a finite effective cosmological constant, recovering the Schwarzschild vacuum in the asymptotic limit. Later developments connected these regular geometries to nonlinear electrodynamics (NLED) [Ayn-Beato1998, Frolov:2016pav], noncommutative geometry [nicolini2009noncommutative] and modified gravity theories [Nojiri2017, NunesdosSantos:2025alw], establishing a broad phenomenological framework in which curvature regularization is associated with an effective matter distribution or quantum corrections of gravitation.Among the physically motivated models of matter sources, the string cloud case, introduced by [letelier1979clouds], stands out. In this approach, the gravitational field is generated by a spherically symmetric distribution of one-dimensional strings radiating from the origin, which modifies the Schwarzschild potential through a constant term proportional to the string parameter $a$. However, Letelier’s solution still presents a curvature singularity at the origin, which prevents a completely self-sustaining physical interpretation.
Recently, several works have sought to regularize the singular structure of Letelier-type geometries using Dymnikova-type exponential regulators [dymnikova1992vacuum, Dymnikova2004]. These regulators smoothly distribute the energy density around a finite scale $r_{0}$, producing spacetimes with de Sitter kernels and finite invariants. However, when combined with the hypergeometric correction that generalizes the original potential of the string cloud [ALENCAR2025102031], the exponential damping becomes insufficient to eliminate all divergences, resulting in residual linear terms in the curvature near the origin. This indicates that standard exponential regularization, while successful for the Schwarzschild and Letelier solutions, does not provide complete regularization when higher-order corrections are incorporated into the string cloud model.
@@
To address these limitations, we introduce a new family of regular black hole solutions sourced by the novel Letelier-Alencar string cloud whose matter distribution is governed by a Dagum-type [dagum2006wealth] rational regulator. This function smooths the energy density near the core, guarantees the finiteness of curvature invariants, and preserves the expected asymptotic behavior. The use of a rational regulator is physically motivated: similar profiles arise in nonlinear electrodynamics models of regular black holes, where they replace the central singularity with a finite-density core [Balart2014]. Here, the Dagum regulator plays the same role, providing a smooth transition between the regular interior and the exterior string-cloud regime.
The resulting spacetime interpolates between a Letelier-dominated exterior and an anti–de Sitter (AdS) vacuum at the origin, whose curvature radius is determined by the regularization scale $r_{0}$ and the string $a$ parameters. The analysis of the parameter space identifies the domains in which horizons form, as well as the critical boundaries where the horizon degenerates, linking macroscopic and microscopic regimes through the interplay between $a$, $r_{0}$, and the mass $M$.
@@
We studied the thermodynamic properties of this regularized configuration in detail, including the dependence of mass, Hawking temperature, entropy, and heat capacity on the model parameters. It is observed that, although $a$ and $r_{0}$ modify the temperature and mass, the entropy depends exclusively on the regularization scale, which reveals a separation between the microscopic degrees of freedom associated with the regular core and those linked to the external string field. Using Rényi entropy formalism, we show that non-additive corrections stabilize the system and smooth the traditional Hawking–Page phase transition, in agreement with recent studies of black holes in non-extensive formalisms [Wei:2022dzw, Okyay_2022]. Furthermore, applying the topological defects approach developed by [wei2022black], we classify the thermodynamic phases by means of the topological charge $W$, demonstrating that non-extensive effects transform the standard two-phase structure (stable and unstable branches) into a single globally stable configuration.
From a phenomenological point of view, the model allows us to look at observational data. By calculating the shadow radius and comparing it with recent measurements of Sgr A* and M87* obtained by the Event Horizon Telescope (EHT) collaboration[akiyama2019first, vgn2025], we establish constraints on the parameters $a$ and $r_{0}$ . The results show that the presence of a regular core tends to reduce the shadow size compared to the Schwarzschild case, a characteristic consistent with other regular geometries [Jiang_2024]. This fact reinforces the physical robustness of the model and suggests that black holes regularized by string clouds can serve as effective descriptions of compact astrophysical objects. In summary, this work presents a regularized black hole solution generated by a generalized string cloud.
@@ III Regular Black Hole Sourced by the New String Cloud
We expand this idea to a new string cloud solution whose gravitational potential includes a hypergeometric correction. This solution was proposed by Alencar et al. in [Alencar:2025zyl]. Here we will work with a more general modified form of the metric function, to be introduced to regularize the geometry while preserving the string cloud contribution, which can initially be written as
@@
where the exponential factor introduces a natural cutoff at scale $r_{0}$, which seeks to ensure regularity near the origin and a smooth transition to the asymptotic region. This scale is made here equal to the characteristic length associated with the string cloud introduced in [Alencar:2025zyl]. The exponential factor thus defines a natural cutoff scale at $r_{0}$ and regulates the spatial profile of the string cloud energy density. The factor containing the exponential term corresponds to a Dymnikova-like mass distribution, which modifies the Schwarzschild geometry to produce regular de Sitter and Schwarzschild vacua in the core and asymptotic regions, respectively. In our framework, for $r\gg r_{0}$, the metric (11) with this Dymnikova-like factor reduces to
@@
The geometry described by Eq. (7), while regular in its metric components, unfortunately leads to divergent curvature invariants in the origin due to the linear term in $r$. This indicates that the Dymnikova-type exponential regulator, while successful for the Schwarzschild and Letelier solutions, is insufficient to fully regularize the spacetime in the presence of the new string cloud. We therefore introduce another regularization scheme based on a rational distribution function, which modifies the metric potential to a new form and successfully ensures the finiteness of all curvature invariants.
@@
which corresponds to an anti–de Sitter vacuum, ensuring the regularity of all curvature invariants at the origin. We note that the chosen Dagum-type regulator represents a smooth, rational function that governs the spatial distribution of the string cloud energy density. Unlike the exponential cutoff, which arises in specific contexts like non-commutative geometry or certain string-theoretic setups, the rational decay of the regularization factor is motivated by its connection to well-established statistical distributions and its proven efficacy in generating regular black holes from nonlinear electrodynamics (NLED) sources [Balart:2014cga]. This regulator effectively “smears” the mass and string tension over a finite region of radius $r_{0}$, mimicking the introduction of a fundamental minimal length scale where classical gravity is modified. The specific choice of the exponents in the rational factor represents the minimal configuration required to keep all curvature invariants finite, yielding a mathematically consistent and physically motivated regular geometry, also inducing a transition to an anti-de Sitter core, a feature often linked to confining vacuum states in gauge/gravity duality.
@@
indicating that the inner region is locally anti-de Sitter with an effective curvature radius $l$ and an effective cosmological constant $\Lambda_{\mathrm{eff}}=-3/l^{2}$. Consequently, both geometries converge toward the same AdS core in the limit $r\to 0$. This correspondence ensures the finiteness of curvature invariants at the origin, guaranteeing the regularity of the spacetime. Deviations between the two metrics appear only at orders equal or greater than $\mathcal{O}(r^{5})$, reflecting the distinct sources of matter involved, the present model being governed by the string cloud and regularization parameters $|a|,r_{0}$, respectively. These higher-order corrections encode the fine structure of the regular core and influence phenomenologically relevant quantities such as the photon-sphere radius and the shadow size.
@@ VI Black holes as topological defects
Recently, Wei et al. [Wei:2022dzw] showed that black holes can be regarded as topological defects. They introduced a off-shell generalized Gibbs free energy
@@
The current (41) is algebraically conserved and vanishes everywhere except at the zeros of $\vec{\phi}$ [Wei:2022dzw]. From Eq. (41), one obtains
@@ VII Shadow analysis
While shadow analyses are often more relevant for rotating geometries, we focus here on the static configuration in order to isolate and systematically evaluate the impact of the string cloud parameters on the shadow morphology. This simplified setup represents a crucial first step toward confronting the present regular black hole model with observational data, which have in turn motivated a number of recent investigations [Liu2019pov, Kimet2021, CIMDIKER2021100900, Okyay_2022, PANTIG2023169197, LIU2024139052, Jiang_2024, vgn2025].
In this section, we estimate the black hole shadow radius and establish preliminary bounds on the free parameters of our model. For computational tractability and to isolate the fundamental geometric effects, we neglect any contribution from accretion flows or plasma and assume a distant observer located in the asymptotic region.

### v2

@@ Abstract
We construct a new family of regular black hole solutions supported by the novel Letelier-Alencar string cloud and regularize it through a rational Dagum-type distribution. The regulator smooths the matter profile and ensures finite curvature invariants, yielding a geometry that interpolates between a string-cloud exterior and an anti-de Sitter core. We analyze the energy conditions, identifying where the null, weak, dominant, and strong conditions hold or fail across the core and exterior, given the regularity of the solution. The parameter space for horizon formation is mapped and the thermodynamic properties – mass, Hawking temperature, entropy, and heat capacity – are derived; notably, the entropy depends only on the regularization scale while the string parameter modifies temperature and heat capacity. Employing Rényi’s non-extensive entropy and the approach of topological thermodynamics, we show that the non-extensive deformation stabilizes the system and eliminates the standard phase transition, leaving only a single thermodynamically stable black hole state. Finally, we compute the shadow radius and derive constraints compatible with the current Event Horizon Telescope bounds for Sgr A* and M87*.
@@ I Introduction
Although the existence of astrophysical BHs at this point is largely taken as a fact due to the incredible fit between models and observations of gravitational waves and BH shadows [LIGOScientific:2016aoc, LIGOScientific:2020ibl, akiyama2019first, EventHorizonTelescope:2022wkp], the formation of singularities is not thought to be as inevitable as it once was [Ashtekar2022, Bambi2023, NOVELLO2008]. The search for regular BH solutions has become one of the most important topics in recent literature, as they can be used to test indirect consequences of quantum and modified gravity theories [Ayn-Beato2005, Fernandes:2020nbq, Ishibashi2021]. Although the singularity theorems developed by Penrose and Hawking indicate that divergences are inevitable under physically reasonable energy conditions, it is widely believed that the fundamental theory of gravitation should remain regular in high curvature regimes, in which quantum effects or nonlinear corrections become relevant.
@@
In this context, the construction of regular black holes (RBHs) offers a particularly useful alternative to eliminate singularities, introducing smooth energy and momentum distributions that produce finite curvature invariants throughout spacetime. Early models, such as those proposed by [Bardeen1973, dymnikova1992vacuum, hayward2006formation], replaced the central singularity with a de Sitter nucleus (dS) characterized by a finite effective cosmological constant, recovering the Schwarzschild vacuum in the asymptotic limit. Later developments connected these regular geometries to nonlinear electrodynamics (NLED) [Ayn-Beato1998, Frolov:2016pav], noncommutative geometry [nicolini2009noncommutative] and modified gravity theories [Nojiri2017, NunesdosSantos:2025alw], establishing a broad phenomenological framework in which curvature regularization is associated with an effective matter distribution or quantum corrections of gravitation. These are only some of the approaches found in the literature to obtain regular BH solutions, but as long as they are fundamentally well motivated, the only way to access their validity is to derive their phenomenological consequences and test their consistency with observations. Broadly speaking, one can try to divide the approaches in three camps: i) solutions obtained from fundamental modifications of gravitational theories, possibly due to quantum effects; ii) solutions obtained from forcing regularity and reconstructing Einstein’s field equations, and iii) solutions obtained from exotic of unusual matter source distributions or non-minimal couplings motivated by fundamental physics.
Among the physically motivated models of matter sources the string cloud case [letelier1979clouds, letelier1981fluids] stands out due to its connection with the possibility that the fundamental entities of nature are string-like which has lead to the historic and still ongoing development of String Theory [Kiritsis2019]. A further motivation comes from the possibility of explaining galactic curves usually associated with Dark Matter (DM) by considering the presence of string fluids with an appropriate equation of state parameter [Soleng:1993yr]. In [Capozziello:2006ij] it was shown that a fluid of strings like the one proposed by Letelier could form part of the DM budget of our universe, as well as offering a Unified Dark Energy (UDE) scenario that could solve the cosmological constant and the coincidence problems [MARTIN2012, Velten2014]. Considering the current challenges facing the standard cosmological model, like the Hubble tension and possible variations of Dark Energy’s equation of state parameter being indicated by DESI, studying further consequences of other matter source models that are fundamentally motivated is of particular importance. In Letelier’s approach, the gravitational field is generated by a spherically symmetric distribution of one-dimensional strings radiating from the origin, which modifies the Schwarzschild potential through a constant term proportional to the string parameter $a$. However, Letelier’s solution still presents a curvature singularity at the origin, which prevents a completely self-sustaining physical interpretation. Recently, several works have sought to regularize the singular structure of Letelier-type geometries using Dymnikova-type exponential regulators [dymnikova1992vacuum, Dymnikova2004, NunesdosSantos:2025alw, Kar:2025phe]. These regulators smoothly distribute the energy density around a finite scale $r_{0}$, producing space-times with de Sitter kernels and finite curvature invariants. Another recent development has been the proposal of a new cloud of string solution that includes a “magnetic-like” contribution to the energy momentum tensor that leads to a modified BH solution with a nonlinear dependence on string length $\ell_{s}$ and string coupling $g_{s}$ coupled to a hypergeometric function of $r^{4}/\ell_{s}^{4}$ [ALENCAR2025102031]. This model provides a more complete physical description beyond Letelier’s initial approach as its dependence on the string coupling can be viewed as encoding string backreaction in the gravitational field at tree level. This solution, which we will refer to as Letelier-Alencar solution, remains singular at the origin and has some limitations when dealing with evaporation up to the scale where quantum backreaction effects should become relevant but serves as an interesting effective extension to the original cloud of strings model and recovers it in the point-particle limit [ALENCAR2025102031]. A natural next step would be to apply the Dymnikova-type regulators that can be viewed as connected to quantum vaccuum corrections [dymnikova1992vacuum, Alencar:2023wyf, Estrada:2023pny, NunesdosSantos:2025alw] and study the regularity of the solutions. However, as we shall demonstrate, when combined with the hypergeometric correction that generalizes the original potential of the string cloud , the exponential damping becomes insufficient to eliminate all divergences, resulting in residual linear terms in the curvature near the origin. This indicates that standard exponential regularization, while successful for the Schwarzschild and Letelier solutions, does not provide complete regularization when higher-order corrections are incorporated into the string cloud model.
@@
To address these limitations, we introduce a new family of regular black hole solutions sourced by the novel Letelier-Alencar string cloud whose matter distribution is governed by a Dagum-type [dagum2006wealth] rational regulator. This function smooths the energy density near the core, guarantees the finiteness of curvature invariants, and preserves the expected asymptotic behavior. The use of a rational regulator is physically motivated: similar profiles arise in nonlinear electrodynamic models of regular black holes, where they replace the central singularity with a finite-density core [Balart2014]. Here, the Dagum regulator plays the same role, providing a smooth transition between the regular interior and the exterior string-cloud regime. The resulting spacetime interpolates between a Letelier-dominated exterior and an anti–de Sitter (AdS) vacuum at the origin, whose curvature radius is determined by the regularization scale $r_{0}$ and the string $a$ parameters. The analysis of the parameter space identifies the domains in which horizons form, as well as the critical boundaries where the horizon degenerates, linking macroscopic and microscopic regimes through the interplay between $a$, $r_{0}$, and the mass $M$.
The most common way to study the link between microscopic and macroscopic aspects of BH solutions is through their thermodynamic behavior. As already mentioned, since the first works of Bardeen and Hawking in the 1970’s the study of BH thermodynamics has become a traditional way of understanding the stability and possible evolution of those objects through evaporation [Bardeen1973, wald2001thermodynamics]. Beyond that, in this context one can study its phase space evolution, discuss phase transitions between different states and consider possible corrections to the standard laws of BH thermodynamics that include thermodynamic pressure from the cosmological constant [Kastor_2009, Rehan2024]. Considering the importance of topology and critical points in the study of classical thermodynamic systems, an approach that has gained a lot of interest in the recent literature consists of using topological charges to understand BH stability, phase transitions and classify solutions in a general manner [Wei:2021vdx, wei2022black]. BH solutions with extended phase spaces present a rich structure of phase transitions, the most widely found transition occurs between small and large solutions in a way that has been compared with the liquid-gas transition of Van de Waals fluids [MANN2012, Wei:2015iwa, Wei:2019uqg, WEI2021]. As the system’s temperature is lowered, a coexistence curve forms in the pressure-temperature plane, delineating the states where small and large black hole phases can exist simultaneously. This curve originates at a critical point and terminates at the origin of the diagram. In the vicinity of this critical point, the system exhibits universal critical behavior, whose properties are accurately described by mean field theory. While the specific shapes of phase diagrams can vary between different systems, the emergence of a critical point is a universal and fundamental feature. We employ this approach to study our regularized Letelier-Alencar string cloud solution and discuss how the two phases emerge.
The analysis of BH thermodynamics can be further enriched by the use of generalized entropies that go beyond the extensive Boltzmann-Gibbs traditional definition [Renyi:1959pbs, Tsallis:1987eu, Tsallis2013]. Beyond the fact that the usual Bekenstein-Hawking formula is explicitly inconsistent with the usual relation $S\propto L^{3}$, modern developments in BH physics indicate that quantum gravitational effects might be connected to entanglement giving further motivation to deal with nonadditive definitions of entropies like the ones proposed by Tsallis or its Renyí-like version [Biro:2013cra, Czinner:2015eyk, Czinner:2017tjq]. It has been shown that for any viable nonadditive entropy, there must exist a transformation—a "formal logarithm" denoted as L(S)L(S)—that converts it into an additive quantity ($L(S_{12})=L(S_{1})+L(S_{2}))$. This allows for a consistent definition of temperature via $1/T=\partial L(S(E))/\partial E$. The Bekenstein-Hawking entropy is a natural example of a nonadditive entropy, following the rule $S_{12}=S_{1}+S_{2}+2\sqrt{(S_{1}S_{2})}$. For more general cases with a non-zero parameter $\lambda$, this framework leads directly to the Rényi entropy. Specifically, when the original entropy follows the Tsallis composition rule $S_{12}=S_{1}+S_{2}+\lambda S_{1}S_{2}$, its formal logarithm is precisely the Rényi entropy: $S_{R}=L(S_{T})$. Both the Tsallis and Rényi formulas generalize standard thermodynamics and reduce to the conventional Boltzmann-Gibbs entropy in the limit where the non-extensivity parameter $q$ approaches 11 $(\lambda\rightarrow 0)$.
@@
We investigate the thermodynamic properties of this regularized configuration in detail, including the dependence of mass, Hawking temperature, entropy, and heat capacity on the model parameters. It is observed that, although $a$ and $r_{0}$ modify the temperature and mass, the entropy depends exclusively on the regularization scale, which reveals a separation between the microscopic degrees of freedom associated with the regular core and those linked to the external string field. Using Rényi entropy formalism, we show that non-additive corrections stabilize the system and smooth the traditional Hawking–Page phase transition, in agreement with recent studies of black holes in non-extensive formalisms [wei2022black, Okyay_2022]. Furthermore, applying the approach of topological defects developed by [wei2022black], we classify the thermodynamic phases by means of the topological charge $W$, demonstrating that non-extensive effects transform the standard two-phase structure (stable and unstable branches) into a single globally stable configuration. From a phenomenological point of view, the model allows us to look at observational data. By calculating the shadow radius and comparing it with recent measurements of Sgr A* and M87* obtained by the Event Horizon Telescope (EHT) collaboration[akiyama2019first, vgn2025], we establish constraints on the parameters $a$ and $r_{0}$. The results show that the presence of a regular core tends to reduce the shadow size compared to the Schwarzschild case, a characteristic consistent with other regular geometries [Jiang_2024]. This fact reinforces the physical robustness of the model and suggests that black holes regularized by string clouds can serve as effective descriptions of compact astrophysical objects. In summary, this work presents a regularized black hole solution generated by a generalized string cloud.
@@ III Regular Black Hole Sourced by the New String Cloud
The Letelier-Alencar string cloud solution provides a more complete framework to study possible consequences of string backreaction into the spacetime structure [ALENCAR2025102031], but it remains a singular solution that might indicate that other effects might have to be taken into account. Here we first try to work with a more general modified form of the metric function, to be introduced to regularize the geometry while preserving the string cloud contribution, which can initially be written as
@@
where the exponential factor introduces a natural cutoff at scale $r_{0}$, which seeks to ensure regularity near the origin and a smooth transition to the asymptotic region. This scale is made here equal to the characteristic length associated with the string cloud introduced in [ALENCAR2025102031]. The exponential factor thus defines a natural cutoff scale at $r_{0}$ and regulates the spatial profile of the energy density of the string cloud. The factor containing the exponential term corresponds to a Dymnikova-like mass distribution (which is connected to quantum vaccum effects [dymnikova1992vacuum]), which modifies the Schwarzschild geometry to produce regular de Sitter and Schwarzschild vaccua in the core and asymptotic regions, respectively. In our framework, for $r\gg r_{0}$, the metric (11) with this Dymnikova-like factor reduces to
@@
The geometry described by Eq. (7), while regular in its metric components, unfortunately leads to divergent curvature invariants in the origin due to the linear term in $r$. This indicates that the Dymnikova-type exponential regulator, while successful for the Schwarzschild and Letelier solutions, is insufficient to fully regularize the spacetime in the presence of the new string cloud. One could in principle consider extending the Dymnikova regulator to a quartic argument in the exponential, but there is no physical justification for such a modification, since higher-order powers are typically associated with higher-dimensional settings [Estrada:2023pny]. We therefore introduce another regularization scheme based on a rational distribution function, which modifies the metric potential to a new form and successfully ensures the finiteness of all curvature invariants.
@@
which corresponds to an anti–de Sitter vacuum, ensuring the regularity of all curvature invariants at the origin. We note that the chosen Dagum-type regulator represents a smooth, rational function that governs the spatial distribution of the string cloud energy density. The rational decay of the regularization factor is motivated by its proven efficacy in generating regular black holes from nonlinear electrodynamics (NLED) sources [Balart:2014cga]. This regulator effectively “smears” the mass and string tension over a finite region of radius $r_{0}$, mimicking the introduction of a fundamental minimal length scale where classical gravity is modified [Frolov:2016pav]. The specific choice of the exponents in the rational factor represents the minimal configuration required to keep all curvature invariants finite, yielding a mathematically consistent and physically motivated regular geometry, also inducing a transition to an anti-de Sitter core, a feature often linked to confining vacuum states in gauge/gravity duality.
@@
indicating that the inner region is locally anti-de Sitter with an effective curvature radius $l$ and an effective cosmological constant $\Lambda_{\mathrm{eff}}=-3/l^{2}$. Consequently, both geometries converge toward the same AdS core in the limit $r\to 0$. Deviations between the two metrics appear only at orders equal or greater than $\mathcal{O}(r^{5})$, reflecting the distinct sources of matter involved, the present model being governed by the string cloud and regularization parameters $a,r_{0}$, respectively. These higher-order corrections encode the fine structure of the regular core and influence phenomenologically relevant quantities such as the photon-sphere radius and the shadow size.
@@ VI Black holes as topological defects
Recently, Wei et al. [wei2022black] showed that black holes can be regarded as topological defects. They introduced a off-shell generalized Gibbs free energy
@@
The current (41) is algebraically conserved and vanishes everywhere except at the zeros of $\vec{\phi}$ [wei2022black]. From Eq. (41), one obtains
@@ VII Shadow analysis
While shadow analyses are often more relevant for rotating geometries, we focus here on the static configuration in order to isolate and systematically evaluate the impact of the string cloud parameters on the shadow morphology. This simplified setup represents a crucial first step toward confronting the present regular black hole model with observational data, which have in turn motivated a number of recent investigations [Liu2019pov, Kimet2021, CIMDIKER2021100900, Okyay_2022, PANTIG2023169197, LIU2024139052, Jiang_2024, vgn2025]. In this section, we estimate the black hole shadow radius and establish preliminary bounds on the free parameters of our model. For computational tractability and to isolate the fundamental geometric effects, we neglect any contribution from accretion flows or plasma and assume a distant observer located in the asymptotic region.